OpenStudy (anonymous):

Solve the system of equations: what is the x-value for the solution? 2x + 7y = 3 and x = 1 - 4y A. -1 B. 5 C. 4 D. 3

OpenStudy (anonymous):

2x + 7y = 3 and x = 1 - 4y 2(1-4y)+7y=3 2-8y+7y=3 2-y=3 y=-1........x=1-4(-1)=5 answer B

OpenStudy (anonymous):

thanks!